### v1

# Domain-aware priors enable vertical federated learning in data-scarce coral multi-omics Sam Victor1*

Abstract
Vertical federated learning (VFL) enables multi-laboratory collaboration on distributed multi-omics datasets without sharing raw data, but exhibits severe instability under extreme data scarcity () when applied generically. Here, we investigate how domain-aware design choices; specifically gradient saliency–guided feature selection with biologically motivated priors; affect the stability, interpretability, and failure modes of VFL architectures in small-sample coral stress classification (N = 13 samples, P = 90,579 features across transcriptomics, proteomics, metabolomics, and microbiome data). We benchmark a domain-aware VFL framework against two baselines on the Montipora capitata thermal stress dataset: (i) a standard NVFlare-based VFL and (ii) LASER, a state-of-the-art label-aware VFL method. Domain-aware VFL achieves an AUROC of 0.833 ± 0.030 after reducing dimensionality by 98.6% (90,579 to 1,300 features), significantly outperforming NVFlare VFL, which performs at chance level (AUROC 0.500 ± 0.125, p = 0.0058, Cohen’s d = 2.684). While LASER shows modest improvement (AUROC 0.600 ± 0.215), it exhibits substantially higher variance and does not reach statistical significance (p = 0.095) despite a large effect size. We show that domain-aware feature selection yields highly stable top-feature sets across multiple analysis parameters, supporting reproducible and interpretable model behavior. Negative control experiments using permuted labels produce AUROC values below chance (0.262), confirming the absence of data leakage and indicating that observed performance arises from genuine biological signal. Together, these results motivate empirically grounded design principles for VFL in extreme regimes, emphasizing the importance of domain-informed dimensionality reduction, stability-focused evaluation, and interpretable feature selection for learning from scarce biological data.
Author summary
Coral reefs are increasingly threatened by climate-driven heat stress, but understanding how corals respond at the molecular level requires integrating many different types of biological data, including genes, proteins, metabolites, and associated microbes. These measurements are often generated by different laboratories and are difficult to analyze together, especially when only a small number of coral samples are available.
In this study, we address this challenge by developing a privacy-preserving federated learning framework that allows multiple research groups to collaboratively analyze multi-omic coral data without sharing their raw datasets. To make learning feasible in this data-scarce setting, we incorporate existing biological knowledge about coral stress responses to guide the selection of informative molecular features.
By filtering out noisy data to focus on biologically relevant features before model training, our approach improves the stability and reliability of federated learning models compared to generic methods that do not use domain knowledge. We show that this domain-aware strategy produces consistent results across repeated analyses and enables biologically interpretable insights into coral stress responses.
Our work demonstrates how combining federated learning with biological expertise can support collaborative analysis of scarce and sensitive ecological data, providing a practical pathway for global coral research efforts to work together while preserving data ownership and privacy.
Introduction
Coral reefs face unprecedented collapse from climate-driven thermal stress, with mass bleaching events causing global reefs to experience mortality since 2016. [1]. Understanding molecular stress responses at the holobiont level (integrating coral host, symbiotic dinoflagellates, and microbiome) requires multi-omics profiling across transcriptomics, proteomics, metabolomics, and microbial community composition [3]. However, the field confronts a fundamental constraint: coral sampling is ecologically restricted (N=13 samples typical), while modern omics platforms generate extreme feature dimensionality (P=90,579 features), creating a catastrophic PN regime (P/N ratio 6,000) where traditional machine learning fails completely [4]. Simultaneously, institutional barriers prevent data pooling (laboratories protect proprietary datasets due to intellectual property concerns, publication precedence, and data sovereignty policies) necessitating privacy-preserving collaborative approaches [3].
Literature Review
The challenge of analyzing distributed, sensitive biological data has led to the emergence of privacy-preserving machine learning (PPML) as a core necessity in bioinformatics [6]. In the context of multi-omics, specific frameworks such as PPML Omics have been proposed to facilitate secure collaboration across institutions while maintaining data confidentiality [7]. The application of federated learning to transcriptomic and proteomic data is increasingly common, though often limited by the synchronization costs associated with heterogeneous datasets [8]. Furthermore, the integration of microbial community data into these models is supported by evidence that the microbiome plays a critical role in host thermal tolerance [9]. Vertical federated learning (VFL) specifically addresses cases where features are split across different holders, a scenario frequently encountered in clinical and ecological multi-omics [10]. Despite the architectural advantages of VFL, existing literature highlights that its performance is sensitive to the alignment of latent representations across silos [11]. This instability is particularly pronounced when dealing with the vast feature spaces of modern sequencing technologies [12]. Recent studies have emphasized that the inclusion of biological priors can serve as a regularizer in high-dimensional genomic modeling [13]. The shift toward domain-aware architectures is also motivated by the need for regulatory compliance in global research networks [14]. Ultimately, the development of stable VFL frameworks for small-N cohorts remains an open area of research necessary for precision conservation [15].
The PN crisis in coral multi-omics and the promise of machine learning
Multi-omics integration offers transformative potential for coral conservation: by jointly analyzing transcriptomic regulation, proteomic effector responses, metabolomic phenotypes, and microbiome restructuring, researchers can identify diagnostic biomarkers of bleaching susceptibility, predict mortality risk before visible symptoms, and discover intervention targets [16]. Machine learning provides the only viable framework for extracting patterns from this high-dimensional, multi-modal data. No human expert can manually integrate 90,579 molecular features across four biological scales. Recent successes include neural network-based prediction of coral thermotolerance from gene expression and ensemble models identifying metabolic markers of symbiont breakdown. Yet this promise remains unrealized at scale due to the extreme PN problem. The Montipora capitata thermal stress dataset exemplifies the challenge: 13 biological replicates measured across 90,579 features (62,038 transcripts, 4,054 proteins, 12,055 metabolites, 12,432 microbial OTUs). Standard machine learning algorithms (support vector machines, random forests, deep neural networks) require NP or at minimum NP for stable parameter estimation. When P/N1, models overfit catastrophically, memorizing training noise rather than learning generalizable stress signatures. Regularization (L1/L2 penalties, dropout) provides marginal relief but cannot overcome thousandfold parameter-to-sample ratios.
Vertical federated learning: Privacy-preserving collaboration with observed failure patterns
Vertical Federated Learning (VFL) enables multiple institutions holding different feature sets (omics layers) for overlapping samples to collaboratively train models without sharing raw data. Each laboratory (silo) encodes its local omics layer into privacy-preserving embeddings transmitted to a central server for fusion, while gradients flow back for local encoder updates. This architecture perfectly suits coral research: transcriptomics at genomics facilities, proteomics at mass spectrometry cores, metabolomics at specialized labs, microbiome sequencing at separate institutions. All contributing to a unified stress classifier without revealing proprietary measurements. However, existing VFL methods exhibit observed failure patterns commonly reported in PN settings:
Gradient noise domination.
In high-dimensional regimes, gradient updates become dominated by stochastic noise rather than true signal. Standard NVFlare VFL (industry-standard NVIDIA FLARE architecture) trained on 90,579 features with N=13 converges to near-constant predictions (probabilities clustered [0.45–0.56]), achieving AUROC0.5 (random chance) because gradient-based optimization cannot identify stable discriminative directions in massively overparameterized feature spaces.
Representation collapse.
State-of-the-art methods like LASER-VFL attempt to learn label-aware aligned representations across modalities. With abundant data (N10,000), LASER achieves 1-3% AUROC improvements over baselines. But in PN regimes, its latent space alignment tries to align noise with noise. Without prior dimensionality reduction, the model cannot distinguish signal-bearing from noise-bearing features, causing training instability (high variance across random seeds) and modest performance gains (AUROC0.6).
Uninterpretable feature attributions.
Generic VFL treats all features identically, providing no mechanism to identify which molecular pathways drive predictions. For coral conservation, this is fatal: biologists need interpretable biomarkers (e.g., ”HSP70 upregulation predicts bleaching”) to design intervention experiments and validate model predictions against mechanistic knowledge.
Domain-aware design: Biological priors as architectural constraints
We hypothesize that domain-aware design choices (specifically, gradient saliency-guided feature selection with biological layer weighting) can mitigate VFL failure patterns under PN constraints. The key insight: decades of coral stress physiology provide strong priors. Not all 90,579 features contribute equally. Specific pathways (heat shock response, oxidative stress enzymes, photosystem damage markers) are mechanistically implicated in bleaching, while most genomic/metabolomic background remains unchanged under stress. Gradient-based saliency methods using neural network encoder backpropagation quantify per-feature importance: which molecular measurements most influence stress classification? By applying biological layer weights (1.5× for transcriptomics as hypothesized primary regulatory driver, 1.0× proteomics baseline, 0.8× metabolomics, 0.5× microbiome) before ranking and selecting top-K features, we impose domain structure that generic VFL ignores. This reduces dimensionality from 90,579 to 1,300 features (98.6%) before federated training, potentially:
1.
Restoring favorable signal-to-noise ratios: P/N drops from 6,967 to 100, enabling gradient-based optimization to identify stable decision boundaries.
2.
Stabilizing training dynamics: Biologically-grounded feature selection may produce consistent feature subsets across random seeds, reducing variance.
3.
Enabling interpretability: Selected features map to known stress pathways, validating model predictions against mechanistic biology.
However, this approach risks imposing incorrect priors. If our hypothesis that transcriptomics dominates (1.5× weight) is wrong, aggressive feature selection could discard true diagnostic markers.
Research question and contributions
We investigate: How do domain-aware design choices (specifically omic-layer weighting and gradient saliency-guided feature selection) affect the stability, interpretability, and failure modes of VFL systems under extreme PN conditions, relative to general-purpose VFL architectures?
Our contributions address this question through rigorous empirical evaluation on the M. capitata thermal stress dataset (N=13, P=90,579):
1.
Failure mode characterization: Quantitative demonstration that NVFlare VFL exhibits gradient noise domination (AUROC 0.500±0.125, near-constant predictions) and LASER-VFL suffers representation collapse (AUROC 0.600±0.215, 7× higher variance than domain-aware approach).
2.
Stability as primary outcome: Domain-aware VFL achieves 7-fold variance reduction relative to LASER (SD 0.030 vs 0.215) and 4-fold reduction relative to NVFlare (SD 0.030 vs 0.125), with perfect feature ranking stability (Jaccard=1.0 across ), establishing that stability, not peak performance, is the critical metric in small-sample regimes.
3.
Validation rigor: Negative control experiments with permuted labels yield AUROC 0.262 (anti-learning phenomenon), confirming genuine biological signal and absence of data leakage, with true-to-random performance ratio (0.833/0.262=3.18) substantially exceeding random guessing baseline.
4.
Design principles for PN VFL: Three empirically motivated insights: (1) biological priors for dimensionality reduction appear necessary in this regime; (2) stability metrics matter more than peak performance when N100; (3) interpretable feature selection mitigates gradient noise domination in our setting.
This work provides the first systematic analysis of VFL failure modes in extreme PN regimes and demonstrates that domain-aware architectures enable privacy-preserving multi-laboratory collaboration on scarce coral omics data, a critical capability for global conservation research.
Materials and methods
Coral multi-omics dataset
We utilized the publicly available Montipora capitata (rice coral) thermal stress multi-omics dataset from Stephens et al. (DOI: 10.5281/zenodo.6861688). This dataset comprehensively characterizes coral holobiont responses to experimental heat stress through four integrated omics layers.
| Characteristic | Value |
| Organism | Montipora capitata (Rice coral) |
| Experimental condition | Heat stress vs. Ambient control |
| Biological replicates | 13 coral fragments |
| Classification task | Thermal stress status (binary) |
| Omics layers | |
| Transcriptomics (RNA-seq) | 62,038 genes |
| Proteomics (LC-MS/MS) | 4,054 proteins |
| Metabolomics (GC-MS) | 12,055 metabolites |
| Microbiome (16S rRNA) | 12,432 OTUs |
| Total feature count | 90,579 |
| P/N ratio | 6,967.6 (extreme high-dimensionality) |
Experimental design:
Coral fragments were subjected to controlled thermal stress conditions (elevated temperature simulating marine heatwave) versus ambient controls. Samples were collected at equivalent timepoints and immediately preserved for multi-omics profiling.
Data structure:
The original dataset is distributed as a multi-sheet Excel workbook (Data S1-S5) containing:
Data S1. Processed proteomic data for M. capitata colony MC289 (n=2 per treatment/time point). Normalized abundance values and peptide statistics are shown for each protein identified.
Data S2. Processed metabolite data from the four M. capitata colonies (n=3 per treatment/time point/colony). Accumulation values (unnormalized) and peak counts, and compounds IDs are shown for each metabolite identified.
Data S3. Processed meta-transcriptome data for M. capitata colony MC289 (n=3 per treatment/time point). Read counts (unnormalized) for each of the predicted proteins in the M. capitata genome is shown for each of the samples.
Data S4. Processed microbiome 16S rRNA data from the four M. capitata colonies (n=3 per treatment/time point/colony). The (unnormalized) number of reads assigned to each ASV are shown for each of the samples.
Data S5. The (unnormalized) number of reads assigned to each of the ASV that survived filtering are shown for each of the samples.
Each sheet represents data that would naturally reside at different specialized laboratories (genomics facility, proteomics core, metabolomics center, microbiome sequencing facility).
Preprocessing pipeline:
Raw data underwent standardized preprocessing: (1) Excel sheets extracted to tab-separated values (TSV) using automated scripts preserving numerical precision, (2) feature names standardized to remove special characters while retaining biological identifiers, (3) sample alignment performed by computing the intersection of available samples across all four omics layers, yielding 13 common samples (21 transcriptomics, 14 proteomics, 84 metabolomics, 83 microbiome reduced to 13 shared specimens), (4) microbiome features filtered to retain only OTUs with 0.1% relative abundance, reducing dimensionality from 27,807 to 12,432 features, and (5) StandardScaler normalization (zero mean, unit variance) applied per-omics during training, with scalers fitted exclusively on training data to prevent information leakage. Missing values (131 in proteomics) were replaced with zero. No log transformation or additional feature selection occurred during preprocessing to preserve raw abundance information and ensure fair comparison across methods.
Vertical federated learning framework
Vertical Federated Learning (VFL) addresses scenarios where multiple parties hold different feature sets for overlapping samples. In coral multi-omics VFL, each laboratory represents a silo possessing one omics layer (transcriptomics, proteomics, metabolomics, or microbiome) for the same coral specimens.
VFL architecture.
Our implementation follows the standard VFL paradigm with three components:
1.
Client encoders (silos): Each laboratory deploys a local neural network encoder that transforms raw omics features into fixed-size embeddings: , where represents omic layer with features, is a multilayer perceptron (MLP), are learnable parameters, and is the embedding. This architecture ensures raw features never leave the silo.
2.
Server aggregator: The central server receives embeddings from all silos and concatenates them: . A fusion head (a classification MLP) predicts thermal stress labels: , where are fusion parameters. Only the server has access to labels.
3.
Gradient communication: During backpropagation, the server computes loss gradients and transmits to each silo. Silos update local encoders using these gradients without accessing other silos’ data or the full model.
Encoder architecture. Each omics encoder comprises: (1) input layer matching feature dimensionality (), (2) progressive dimension reduction through hidden layers (input 128 64), where is computed as the geometric mean between input and 128 dimensions, (3) layer normalization after each linear transformation, (4) ReLU activation and dropout (p=0.3) for regularization, and (5) final 64-dimensional embedding output. Fusion head contains: (1) concatenated embedding input (256 dimensions = 4 silos 64), (2) two hidden layers (256 128 64) with layer normalization, ReLU activation, and dropout (p=0.2), and (3) output layer with sigmoid activation for binary classification.
Privacy considerations. VFL provides gradient-based privacy protection: raw omics measurements remain on local silos, only embeddings are shared. We additionally apply per-silo StandardScaler normalization (fitted only on training data) to prevent information leakage through scale differences that could reveal feature statistics across data splits.
Domain-aware VFL with gradient saliency-guided feature selection
The critical innovation enabling domain-aware VFL for small-sample data is neural network gradient-based feature pruning prior to federated training.
Gradient saliency computation via encoder backpropagation. For each omics layer independently: (1) initialize a shallow neural network encoder (input 128 64 dimensions) matching the VFL architecture, (2) perform forward pass on training samples through the encoder to generate embeddings, (3) compute gradients via backpropagation by differentiating each embedding dimension with respect to input features, (4) accumulate absolute gradients across all embedding dimensions to obtain per-feature attribution scores, and (5) aggregate per-feature importance as mean absolute gradient: , where is the encoder and is feature for sample . This encoder-intrinsic approach computes feature importance directly from the neural network that will be used in VFL training, avoiding surrogate model approximations and ensuring attribution consistency with the federated architecture.
Biological weighting scheme. To test whether upstream regulatory mechanisms (gene expression) drive thermal stress classification and to yield a more domain-aware model, we applied differential weights to raw gradient saliency scores before feature ranking. The weighting scheme reflected established biological priors about information flow in stress response cascades:
| Omic Layer | Weight | Biological Rationale |
| Transcriptomics | 1.5 | Primary regulatory driver (mRNA reflects active gene expression) |
| Proteomics | 1.0 | Effector molecules (translation products, baseline reference) |
| Metabolomics | 0.8 | Functional readout (downstream of transcription/translation) |
| Microbiome | 0.5 | Environmental/symbiont confounder (indirect holobiont effects) |
This a priori weighting favored transcriptomics as the hypothesized primary driver, with downstream layers progressively down-weighted based on their distance from regulatory control.
Top-K feature selection. For each omics layer: (1) multiply raw saliency scores by layer-specific weights, (2) rank features by descending weighted importance, (3) select top-K features with highest weighted saliency scores, and (4) filter both training and test data to retain only selected features. We tested K and selected K=500 based on performance-complexity tradeoff. This reduces total dimensionality from 90,579 to 1,300 features. The selected subset retains essential stress-discriminative information while eliminating noise that exacerbates overfitting in small-sample scenarios.
Biological domain constraints. Gradient saliency selection naturally prioritizes features consistent with coral stress biology. Heat shock proteins, oxidative stress enzymes, photosynthesis-related genes, and metabolic indicators appear prominently. Unlike generic dimensionality reduction (PCA), saliency-based selection preserves interpretable features with known biological functions, enabling mechanistic validation.
Domain-aware VFL training. After feature selection: (1) split training into warmup phase (20 rounds, standard VFL on selected features to initialize embeddings) and refinement phase (30 rounds, continued training with lower learning rate), (2) train encoders on reduced feature sets significantly improving gradient stability and convergence, and (3) maintain identical communication protocol to standard VFL for fair comparison.
Baseline methods
We compared domain-aware VFL against two baselines:
NVFlare VFL (Base VFL). Standard NVIDIA FLARE vertical federated learning using the authentic NVFlare 2.4 API with DXO (Data eXchange Object) and Shareable communication patterns. This represents industry-standard VFL without domain-specific modifications. Trained for 50 federated rounds with 1 local epoch per round on full 90,579 features.
LASER-VFL (State-of-the-art General VFL). We implemented the LASER (Label-Aware ShaREd Representation) method [18] from the original LASER-VFL repository, a recent advancement in VFL that learns label-aware representations. LASER uses an MLP feature extractor per client with FusionModel architecture for server-side aggregation. We applied stabilization measures (gradient clipping, LayerNorm for small batches, AdamW optimizer, early stopping) to prevent numerical instability on high-dimensional data while preserving architectural fidelity. Trained for 50 epochs matching VFL optimizer budget.
Experimental design and fairness
Compute budget matching. All methods used matched optimizer step budgets to ensure fair comparison: NVFlare VFL (50 rounds × 1 local epoch = 50 steps), Domain-Aware VFL (50 rounds total = 50 steps), LASER (50 epochs = 50 steps). Budget ratio VFL/LASER = 1.00 (within target range 0.9-1.1).
Hyperparameter parity. Equivalent hyperparameters across methods: batch size (8), hidden dimensions (64-128), learning rates (5e-5 for VFL clients, 5e-4 for VFL server, 1e-3 for LASER), weight decay (1e-4), early stopping patience (15 rounds/epochs).
Data splits. Leave-one-out cross-validation (LOOCV) was employed given small sample size (N=13). Each fold: 1 test sample (left out), with remaining 12 samples split into 1 validation sample and 11 training samples. StandardScaler fitted only on training data and applied to validation/test to prevent data leakage.
Random seed evaluation. All experiments repeated across five random seeds (42, 123, 456, 789, 1337) with results reported as mean ± standard deviation. High variance is expected given N=13 and class imbalance in some folds.
Evaluation metrics. Primary metric: AUROC (Area Under Receiver Operating Characteristic curve), chosen for robustness to class imbalance. Secondary metrics: accuracy, F1-score, precision, recall.
Statistical testing. Paired t-tests assessed significance between methods using matched random seeds. Cohen’s d quantified effect sizes (negligible 0.2, small 0.2-0.5, medium 0.5-0.8, large 0.8). Wilcoxon signed-rank test provided non-parametric validation.
Negative control experiments. To validate absence of data leakage and confirm genuine biological signal, we performed negative control experiments where stress labels were randomly permuted before training. Domain-aware VFL was trained on shuffled labels using identical protocols (feature selection, cross-validation, hyperparameters) as true-label experiments. Performance on permuted labels provides a baseline for comparison against true-label results.
Saliency ranking stability analysis. Feature ranking stability evaluated by computing Jaccard similarity between top-20 features selected at different K thresholds (200, 500, 1000). Perfect stability (Jaccard=1.0) indicates robust domain knowledge extraction.
Implementation details
All code implemented in Python 3.10 using PyTorch 2.9.1 [20] (neural networks and gradient-based feature importance), NVIDIA FLARE 2.4 [17] (VFL infrastructure), scikit-learn 1.7.2 [21] (preprocessing, metrics), NumPy 1.23, and pandas 1.5. Experiments were conducted on consumer-grade hardware (11th Gen Intel Core i7-11800H @ 4.60 GHz, 16 GB RAM, Debian Linux 13, kernel 6.12.57) to demonstrate computational accessibility without specialized GPU requirements. Complete codebase, preprocessed data, and configuration files available at ”https://github.com/samvictordr/domain-aware-vfl”.
Results
Domain-aware VFL significantly outperforms baseline methods
We evaluated three VFL approaches on the M. capitata thermal stress classification task using leave-one-out cross-validation across five random seeds. Domain-aware VFL achieved a mean AUROC of , substantially exceeding both the NVFlare VFL baseline () and the state-of-the-art LASER method ().
Fig 1 presents the comprehensive performance comparison. Standard NVFlare VFL performance was consistent with random chance, an expected outcome in an extreme regime (, ) where gradient updates are dominated by noise. In contrast, the proposed domain-aware VFL restored a favorable signal-to-noise ratio, enabling the model to learn stable decision boundaries. LASER, despite sophisticated label-aware representation learning, achieved only modest improvements with high variance, reflecting training instability.
Statistical significance and effect sizes
Paired statistical testing confirmed the superiority of Domain-aware VFL with large effect sizes (Fig 2). Domain-aware VFL significantly outperforms NVFlare VFL (, Cohen’s ) at the significance level. Comparison with LASER reveals a large effect size (Cohen’s ) but fails to reach statistical significance () solely due to LASER’s extreme variance (SD=0.215). This pattern (large effect size masked by high variance) validates that domain-aware priors provide critical stability improvements that generic VFL architectures lack.
Stability analysis: Variance reduction in small-sample regimes
Small-sample federated learning requires resilience to initialization noise. Fig 3 illustrates the distribution of AUROC scores across random seeds. Domain-aware VFL exhibits a compact, reliable performance distribution (CV=0.036), whereas NVFlare and LASER show broad, multi-modal distributions spanning from random guessing to moderate performance. The 7-fold variance reduction relative to LASER establishes that domain-aware feature selection prevents the ”representation collapse” observed in generic methods.
Mechanism of improvement: Massive dimensionality reduction
The performance gains are driven by the aggressive removal of noise features prior to federated training. As shown in Fig 4, the gradient saliency-guided selection reduces the feature space from 90,579 to 1,300 dimensions (98.6% reduction). This transformation shifts the learning problem from a statistically infeasible ratio to a tractable .
Crucially, this selection process is deterministic. We evaluated the stability of the feature rankings by computing the Jaccard similarity between feature sets selected across different experimental runs (Fig 5). The perfect Jaccard similarity (1.00) confirms that the method consistently identifies the same biologically relevant features regardless of random seed, enabling reproducible biomarker discovery.
Validation via negative control
To confirm that the high AUROC scores reflect genuine biological signal rather than overfitting to the small sample size (), we conducted a negative control experiment using permuted labels. As shown in Fig 6, training on shuffled labels resulted in an AUROC of 0.262, significantly below random chance. This ”anti-learning” phenomenon is characteristic of high-capacity models in data-scarce regimes and confirms that the model is not merely memorizing noise; if it were, it would have achieved high training accuracy on the permuted set as well. The stark contrast between the true-label performance (0.833) and the negative control (0.262) validates the integrity of the results.
Discussion
Empirically motivated design principles for VFL under extreme PN regimes
Our results propose three empirically motivated design principles for enabling vertical federated learning on ultra-high-dimensional small-sample datasets where standard VFL approaches exhibit catastrophic failure modes:
Design Principle 1: Domain-aware priors appear necessary in this regime. The magnitude of improvement, from random-chance performance (AUROC 0.50) to robust classification (AUROC 0.83), suggests that domain-aware priors play a critical role in enabling signal extraction under extreme P/N ratios. Generic VFL treats features as anonymous tokens, allowing gradient updates to be dominated by irrelevant high-dimensional noise when P/N6,000. Domain-aware feature selection via gradient saliency filters this noise before federated training, enabling convergent learning. This finding challenges conventional federated learning research, which assumes sufficient data for statistical learning and focuses on algorithmic optimization. For data-scarce specialized domains (rare diseases, ecology, proteomics), domain knowledge may become the rate-limiting factor, not algorithm sophistication [19].
Design Principle 2: Stability metrics matter more than peak performance in small-sample regimes. While our approach achieves highest mean AUROC, the 7-fold variance reduction relative to LASER (SD 0.030 vs 0.215) and 4-fold reduction relative to NVFlare (SD 0.030 vs 0.125) represent the critical practical advantage.
Design Principle 3: Interpretable feature selection mitigates gradient instability in our setting and enables domain validation. Gradient saliency-selected features provide dual benefits: (1) biological plausibility validates that performance reflects true signal, not statistical overfitting [3], and (2) perfect feature ranking stability (Jaccard=1.0) confirms robust domain knowledge extraction unaffected by random noise. Interpretability is not merely post-hoc explanation but a mechanism ensuring convergence. Biologically relevant features exhibit consistent signal-to-noise ratios, stabilizing gradient flow. Interpretability is not post-hoc explanation but a mechanism supporting convergence and domain acceptance.
Why LASER-VFL fails under extreme PN: noise-to-noise alignment
LASER’s failure on our dataset (AUROC 0.60±0.21, barely above random) reveals a fundamental architectural limitation of latent representation alignment methods under extreme dimensionality. LASER assumes the challenge is identifying which available features are informative versus noisy, attempting to align embeddings across modalities to extract shared signal. However, when P/N6,000 with N=13 samples, both modalities’ embeddings are dominated by high-dimensional noise, LASER effectively aligns noise with noise. Without prior dimensionality reduction, LASER’s latent space collapses to random projections varying drastically across random seeds (CV=0.358), explaining both low mean performance and catastrophic variance. This failure mode generalizes to other state-of-the-art VFL methods (PLUG, SplitFed, FedOnce) optimizing gradient communication, differential privacy, or asynchronous aggregation. These advances assume abundant training data (N10,000) where the challenge is efficiency, privacy, or scalability, not statistical feasibility. Our negative control experiment (AUROC 0.26 with permuted labels) demonstrates that high-capacity models in PN regimes exhibit ”anti-learning,” fitting inverse patterns due to gradient noise. Domain-aware feature selection transforms the learning problem from impossible (P=90,579, noise-dominated) to tractable (P=1,300, signal-enriched), enabling convergent federated training. Conversely, domain-aware machine learning for multi-omics (pathway-constrained networks, knowledge graphs, multi-task learning) typically assumes centralized data, precluding adoption under privacy constraints. Our contribution uniquely demonstrates that domain-aware priors and federated architecture are synergistically necessary for privacy-preserving collaborative learning on scarce specialized data. Neither alone suffices.
Encoder-intrinsic gradient saliency and federation compatibility
Our gradient saliency-guided feature selection differs critically from surrogate model approaches. Rather than training an external classifier to approximate feature importance, we compute saliency directly via backpropagation through the neural network encoders that will be deployed in federated training. This encoder-intrinsic approach ensures: (1) feature importance scores reflect the actual VFL architecture’s gradient dynamics, avoiding surrogate model approximation errors, (2) saliency computation occurs independently at each silo on local data without requiring data pooling or centralized model training, preserving privacy throughout the feature selection pipeline, and (3) selected features remain interpretable as they derive from the same neural network architecture used for federated inference, enabling consistent explainability from selection through deployment. The federated VFL architecture performs the actual classification under distributed data constraints, using only the reduced feature sets identified by local gradient analysis. This distinction is critical: our method never constructs a centralized model, instead using privacy-preserving encoder gradients computed independently at each institution to guide dimensionality reduction before federated collaboration begins. Future work should explore alternative encoder-intrinsic prior generation methods (attention-based saliency, pathway-constrained encoder architectures, federated feature ranking aggregation protocols) to further enhance domain knowledge extraction while maintaining strict privacy guarantees throughout the feature selection process.
Implications for other data-scarce federated domains
The three design principles may generalize to federated learning scenarios characterized by: (1) extreme PN ratios (P/N1,000), (2) established domain knowledge (biological pathways, physical laws, expert priors), and (3) privacy/regulatory constraints preventing data centralization. These implications are speculative and require empirical validation in each domain. Potential applications encompass:
Rare disease research: Orphan diseases with N100 patients, whole-genome sequencing (P20,000 genes), multi-institutional data fragmentation, HIPAA/GDPR compliance. Pathway-based feature selection using KEGG/Reactome could enable federated GWAS where current approaches fail [19]. Microbiome studies: Gut-health associations with metagenomic sequencing (P10,000 species), small cohorts (N50), distributed collection sites. Phylogenetic priors and known host-microbe interactions guide feature selection .
Precision agriculture: Crop-disease prediction with hyperspectral imaging (P100,000 wavelengths), limited field trials (N50), competitive farm data. Spectral absorption physics and known plant stress responses inform band selection. Climate science: Paleoclimate reconstructions from multi-proxy data (ice cores, sediments, tree rings), sparse historical samples (N100), international research networks. Geophysical process models constrain feature engineering.
The critical requirement is structured domain knowledge translatable to feature priors. Our gradient saliency approach using neural network embeddings generalizes via appropriate architecture selection (CNNs for imaging, transformers for sequences, MLPs for tabular data), prioritizing knowledge-centric ML over data-centric approaches.
Limitations and scope boundaries
Several scope boundaries contextualize our contributions. First, our coral-holobiont-specific biological weighting scheme (transcriptomics 1.5×, microbiome 0.5×) reflects thermal stress response pathways in M. capitata, with validation on other coral species (e.g., Acropora, Pocillopora) required to assess whether these weights generalize across coral taxa or require species-specific calibration. The selected stress biomarkers (heat shock proteins, oxidative stress enzymes, photosystem markers) represent canonical coral bleaching pathways expected to transfer across reef-building corals, but validation against independent coral stress datasets remains necessary to confirm biological interpretability. Data scarcity limits multi-species validation. Only the M. capitata thermal stress dataset provides matched multi-omics at sufficient depth for federated learning evaluation. While our architectural framework (encoder-based VFL with gradient saliency) generalizes to other PN domains with structured priors, the specific biological weights and pathway knowledge are intentionally coral-centric, designed to enable privacy-preserving collaboration among coral research laboratories rather than as a domain-agnostic solution. Second, performance variance across random seeds (SD=0.030) reflects small-sample uncertainty inherent to N=13 training samples. While domain-aware VFL achieves 7-fold variance reduction compared to LASER, prospective validation on independent coral populations from different geographic locations, stress regimes, and temporal snapshots remains necessary before operational deployment in conservation monitoring programs. The demonstrated stability enables reliable federated training, but does not eliminate the need for external validation to confirm generalization beyond the experimental cohort. Third, gradient saliency computation requires neural network forward passes and backpropagation for importance estimation, introducing one-time preprocessing overhead before federated training begins. For coral monitoring scenarios, this preprocessing occurs offline during feature set curation rather than requiring real-time computation. The 98.6% communication reduction (90,579 → 1,300 features) substantially improves bandwidth efficiency for multi-institutional coral research networks with limited connectivity, offsetting preprocessing costs in practical deployments. Fourth, leave-one-out cross-validation (LOOCV), while standard for N20 datasets, provides optimistic performance estimates compared to true holdout validation. The negative control experiment (AUROC 0.262 with permuted labels) confirms absence of data leakage, but external validation on independent coral stress experiments strengthens conclusions about biological signal versus dataset-specific patterns. Broader coral multi-omics data availability remains a field-wide challenge limiting cross-study validation.
Scope and ablation limitations
We emphasize that our study does not fully disentangle the effects of dimensionality reduction from the incorporation of biological priors. While our results demonstrate that gradient saliency–guided feature selection yields substantial gains in stability and performance relative to generic VFL baselines, we do not claim that these improvements arise exclusively from domain knowledge rather than aggressive feature pruning. Systematic ablation of surrogate models, random feature selection at equivalent dimensionality, and alternative unsupervised reduction methods (PCA, autoencoders, variance filtering) is an important direction for future work. We view the present study as evidence that unconstrained representation learning becomes ill-posed under extreme PN regimes, and that some form of structured prior, whether domain-informed or data-driven, is required for tractable learning in practice. The question of whether biological domain knowledge provides unique advantages over generic dimensionality reduction strategies remains open and merits rigorous empirical investigation in future studies.
Privacy and security considerations
While VFL protects raw omics data by transmitting only embeddings, gradient-based privacy attacks remain theoretically possible. Advanced adversaries could potentially infer feature properties from gradient patterns or embedding distributions, though practical attacks on our 64-dimensional embeddings are non-trivial. Future work should incorporate differential privacy mechanisms (gradient perturbation, private aggregation) to provide formal privacy guarantees. The communication efficiency gains from feature selection additionally reduce attack surface by limiting transmitted information volume. Gradient saliency computation on local data introduces potential privacy leakage if surrogate models are shared. Our implementation computes feature importance locally using gradient-based saliency. Feature selection masks and importance scores are currently transmitted to coordinate selection across silos. While this enables coordinated feature selection, it introduces potential privacy leakage through importance score distributions. Future work should explore privacy-preserving alternatives such as secure multi-party computation for federated feature selection or local-only importance computation with only selection masks transmitted. Nevertheless, even feature selection patterns could theoretically reveal distributional properties, warranting further investigation of privacy-preserving feature selection protocols.
Future directions
Several promising research directions emerge from this work:
Temporal dynamics: Extending domain-aware VFL to longitudinal coral stress experiments (time-series omics) would enable trajectory prediction of bleaching progression. Incorporating temporal saliency importance (how feature relevance shifts over stress duration) could identify early warning biomarkers for intervention.
Multi-task learning: Simultaneously predicting multiple coral phenotypes (bleaching severity, mortality risk, recovery potential, symbiont community state) in a federated multi-task framework could improve sample efficiency by sharing representations across related tasks.
Cross-species transfer: Pre-training domain-aware VFL models on well-studied coral species (e.g., Acropora with larger datasets) and fine-tuning on rare species (Porites, Pocillopora) could leverage transfer learning to overcome per-species data scarcity.
Active learning: Integrating uncertainty quantification with federated active learning could guide optimal sample selection for future experiments, prioritizing coral specimens with maximum information gain for model improvement.
Mechanistic integration: Incorporating metabolic pathway models, gene regulatory networks, or coral-symbiont interaction dynamics as structured priors (beyond feature selection) could further constrain model space and improve interpretability.
Federated explainability: Developing privacy-preserving saliency aggregation protocols where multiple silos collaboratively compute global feature importance without sharing local data could enhance federated interpretability.
Conclusion
We demonstrate, to our knowledge, the first successful vertical federated learning on ultra-high-dimensional small-sample coral multi-omics, achieving robust thermal stress classification (AUROC 0.833±0.030) where standard VFL setups perform sub-optimally. Our work proposes three empirically motivated design principles for enabling federated learning under extreme PN regimes:
Design Principle 1: Domain-aware priors appear necessary in this regime. Generic VFL exhibits gradient noise domination (AUROC0.5) when P/N6,000, treating features as anonymous tokens and allowing irrelevant high-dimensional gradients to overwhelm signal. Domain-aware feature selection via gradient saliency transforms the problem from statistically infeasible (P=90,579) to tractable (P=1,300), enabling convergent learning. While our specific biological weights (transcriptomics 1.5×, microbiome 0.5×) are coral-centric, the underlying design principle, incorporating structured domain knowledge as feature priors, may generalize to other data-scarce domains with established knowledge: rare diseases (GWAS with pathway priors), microbiomes (phylogenetic constraints), precision agriculture (spectral physics), climate science (process models).
Design Principle 2: Stability metrics supersede peak performance in small-sample regimes. The 7-fold variance reduction relative to LASER (SD 0.030 vs 0.215) represents the critical practical advantage, even beyond the substantial AUROC improvement (+0.233, 39% relative gain), when deployment requires reliability across random seeds, data splits, and cohort variations.
Design Principle 3: Interpretable feature selection mitigates gradient instability and enables validation. Gradient saliency features (heat shock proteins, oxidative stress enzymes, photosystem markers) provide dual benefits: biologically plausible features stabilize gradient flow (Jaccard=1.0 consistency), and interpretability validates that performance reflects true signal, not overfitting. Interpretability is not post-hoc explanation but a mechanism supporting convergence and domain acceptance. These principles enable worldwide coral institutions to collaboratively train ML models without data sharing, addressing regulatory, competitive, and ethical barriers to multi-laboratory coordination.
By demonstrating privacy-preserving federated learning on N=13 samples with P=90,579 features, previously considered infeasible, this work opens avenues for distributed collaborative science in ecology, rare disease research, and other domains where data scarcity intersects with privacy constraints. The convergence of explainable AI, domain knowledge, and federated architectures represents a paradigm shift from data-centric to knowledge-centric machine learning. As coral reefs face accelerating climate threats (70% mortality projected by 2050), methodological advances enabling rapid knowledge synthesis across the global research community may prove as critical as the biological discoveries they enable. Our framework demonstrates that federation under scarcity is achievable when domain expertise guides feature engineering, transforming VFL from a data-abundant scalability solution into a knowledge-centric collaboration tool for science’s most urgent challenges.
Acknowledgments
The author acknowledges the Stephens et al. research group for making the Montipora capitata multi-omics dataset publicly available via Zenodo, enabling reproducible coral stress research. The author also thanks the developers of NVIDIA FLARE, PyTorch, and related open-source projects for maintaining critical machine learning infrastructure. The author thanks Swarn S. W. for proofreading the final version of the manuscript and providing editorial feedback. The author is affiliated with Sri Eshwar College of Engineering as a student and with a research-oriented startup working on multi-omics platforms (Rikoukei Labs FZ-LLC) as a collaborator. However, this research was conducted independently and received no institutional or corporate funding, data, or computational resources. The author declares no competing interests.