### v3

# Domain-aware priors stabilize, not merely enable, vertical federated learning in data-scarce coral multi-omics Sam Victor1*

Abstract
Vertical federated learning (VFL) enables multi-laboratory collaboration on distributed multi-omics datasets without sharing raw data, but exhibits severe instability under extreme data scarcity () when applied generically. Here, we investigate how domain-aware design choices—specifically gradient saliency-guided feature selection with biologically motivated priors—affect the stability, interpretability, and failure modes of VFL architectures in small-sample coral stress classification (N = 13 samples, P = 90,579 features across transcriptomics, proteomics, metabolomics, and microbiome data). We benchmark REEF, a domain-aware VFL framework, against two baselines on the Montipora capitata thermal stress dataset: (i) a standard NVFlare-based VFL and (ii) LASER, a state-of-the-art label-aware VFL method. REEF achieves an AUROC of 0.776 ± 0.039 after reducing dimensionality by 98.6% (90,579 to 1,300 features), substantially outperforming NVFlare VFL at chance level (AUROC 0.500 ± 0.125, , Cohen’s ) and numerically exceeding LASER (AUROC 0.557 ± 0.191, , Cohen’s ), with 3–5-fold variance reduction (SD 0.039 vs 0.125 and 0.191). An equal-weights ablation holding feature count constant at 1,300 but removing domain priors confirms that biological priors specifically contribute stability: removing priors yields statistically indistinguishable mean AUROC () but 2.3 higher variance (CV 0.110 vs 0.050). Domain-aware feature selection yields deterministic, budget-invariant feature rankings (Jaccard=1.0 across selection thresholds). Negative control experiments using permuted labels produce AUROC near or below chance (0.357 for REEF, 0.238 for NVFlare VFL), consistent with the absence of gross data leakage; however, with , this control rules out inflated performance rather than formally confirming below-chance behavior. These results motivate design principles for VFL in extreme regimes, emphasizing domain-informed dimensionality reduction, stability-focused evaluation, and interpretable feature selection for scarce biological data.
Author summary
Coral reefs are increasingly threatened by climate-driven heat stress, but understanding how corals respond at the molecular level requires integrating multiple types of biological data, such as genes, proteins, metabolites, and microbes. These measurements are often generated by separate laboratories and difficult to combine, especially when only a small number of coral samples are available.
In this study, we address this challenge by developing a privacy-preserving federated learning framework enabling multiple research groups to collaboratively analyze multi-omic coral data without sharing raw data. To address data scarcity, we incorporate biological knowledge about coral stress responses to guide the selection of informative molecular features.
By filtering out noisy data to focus on biologically relevant features before model training, our approach improves the stability and reliability of federated learning models compared to generic methods that do not use domain knowledge. We show that this domain-aware strategy produces consistent results across repeated analyses and enables biologically interpretable insights into coral stress responses.
Our work demonstrates how combining federated learning with biological expertise can support collaborative analysis of scarce and sensitive ecological data, providing a practical pathway for global coral research efforts to work together while preserving data ownership and privacy.
Introduction
Coral reefs face unprecedented collapse from climate-driven thermal stress, with mass bleaching events causing global reefs to experience mortality since 2016 [1]. Understanding molecular stress responses at the holobiont level (integrating coral host, symbiotic dinoflagellates, and microbiome) requires multi-omics profiling across transcriptomics, proteomics, metabolomics, and microbial community composition [3]. However, the field confronts a fundamental constraint: coral sampling is ecologically restricted (N=13 samples typical), while modern omics platforms generate extreme feature dimensionality (P=90,579 features), creating a severe PN regime (P/N ratio 6,000) where traditional machine learning fails completely [4]. Simultaneously, institutional barriers prevent data pooling (laboratories protect proprietary datasets due to intellectual property concerns, publication precedence, and data sovereignty policies) necessitating privacy-preserving collaborative approaches [3].
Literature review
The challenge of analyzing distributed, sensitive biological data has led to the emergence of privacy-preserving machine learning (PPML) as a core necessity in bioinformatics. In the context of multi-omics, specific frameworks such as PPML Omics have been proposed to facilitate secure collaboration across institutions while maintaining data confidentiality [6, 7]. Centralized graph-based approaches such as MOGONET [5] have demonstrated the potential of multi-omics integration for biomedical classification, but these methods require pooling raw data and thus cannot satisfy institutional privacy constraints. Furthermore, the integration of microbial community data into these models is supported by evidence that the microbiome plays a critical role in host thermal tolerance [8]. Vertical federated learning (VFL) specifically addresses cases where features are split across different holders, a scenario frequently encountered in clinical and ecological multi-omics [9]. Despite the architectural advantages of VFL, existing literature highlights that its performance is sensitive to the alignment of latent representations across silos [10]. Recent studies have emphasized that the inclusion of biological priors can serve as a regularizer in high-dimensional genomic modeling [11]. Ultimately, the development of stable VFL frameworks for small-N cohorts remains an open area of research necessary for precision conservation [13].
The PN crisis in coral multi-omics and the promise of machine learning
Multi-omics integration offers transformative potential for coral conservation: by jointly analyzing transcriptomic regulation, proteomic effector responses, metabolomic phenotypes, and microbiome restructuring, researchers can identify diagnostic biomarkers of bleaching susceptibility, predict mortality risk before visible symptoms, and discover intervention targets [14]. The Montipora capitata thermal stress dataset exemplifies the challenge: 13 biological replicates measured across 90,579 features (62,038 transcripts, 4,054 proteins, 12,055 metabolites, 12,432 microbial OTUs). Standard machine learning algorithms (support vector machines, random forests, deep neural networks) require NP or at minimum NP for stable parameter estimation. When P/N1, models overfit severely, memorizing training noise rather than learning generalizable stress signatures.
Vertical federated learning: privacy-preserving collaboration with observed failure patterns
Vertical Federated Learning (VFL) enables multiple institutions holding different feature sets (omics layers) for overlapping samples to collaboratively train models without sharing raw data. Each laboratory (silo) encodes its local omics layer into privacy-preserving embeddings transmitted to a central server for fusion, while gradients flow back for local encoder updates. This architecture perfectly suits coral research: transcriptomics at genomics facilities, proteomics at mass spectrometry cores, metabolomics at specialized labs, microbiome sequencing at separate institutions—all contributing to a unified stress classifier without revealing proprietary measurements. However, existing VFL methods exhibit observed failure patterns commonly reported in PN settings:
Gradient noise domination.
In high-dimensional regimes, gradient updates become dominated by stochastic noise rather than true signal. Standard NVFlare VFL (industry-standard NVIDIA FLARE architecture) trained on 90,579 features with N=13 converges to near-constant predictions (probabilities clustered [0.45–0.56]), achieving AUROC0.5 (random chance) because gradient-based optimization cannot identify stable discriminative directions in massively overparameterized feature spaces.
Representation collapse.
State-of-the-art methods like LASER-VFL attempt to learn label-aware aligned representations across modalities. With abundant data (N10,000), LASER achieves 1-3% AUROC improvements over baselines. But in PN regimes, its latent space alignment tries to align noise with noise. Without prior dimensionality reduction, the model cannot distinguish signal-bearing from noise-bearing features, causing training instability (high variance across random seeds) and modest performance gains (AUROC0.56).
Uninterpretable feature attributions.
Generic VFL treats all features identically, providing no mechanism to identify which molecular pathways drive predictions. For coral conservation, this is fatal: biologists need interpretable biomarkers (e.g., ”HSP70 upregulation predicts bleaching”) to design intervention experiments and validate model predictions against mechanistic knowledge.
Domain-aware design: biological priors as architectural constraints
We hypothesize that domain-aware design choices (specifically, gradient saliency-guided feature selection with biological layer weighting) can mitigate VFL failure patterns under PN constraints. The key insight: decades of coral stress physiology provide strong priors. Not all 90,579 features contribute equally. Specific pathways (heat shock response, oxidative stress enzymes, photosystem damage markers) are mechanistically implicated in bleaching, while most genomic/metabolomic background remains unchanged under stress. Gradient-based saliency methods using neural network encoder backpropagation quantify per-feature importance: which molecular measurements most influence stress classification? We impose domain structure that generic VFL ignores through two mechanisms: omic-layer feature budgets (controlling how many features are selected per layer) and biological weights applied to encoder embeddings during training (1.5× transcriptomics, 1.0× proteomics, 0.8× metabolomics, 0.5× microbiome), reflecting hypothesized regulatory importance in the stress response cascade. This reduces dimensionality from 90,579 to 1,300 features (98.6%) before federated training, with three hypothesized benefits: restoring a tractable P/N ratio (6,967 → 100), stabilizing gradient updates by selecting biologically consistent feature subsets, and enabling interpretability by retaining features with known mechanistic roles in coral stress pathways.
However, this approach risks imposing incorrect priors. If our hypothesis that transcriptomics dominates (1.5× weight) is wrong, aggressive feature selection could discard true diagnostic markers.
Research question and contributions
We investigate: How do domain-aware design choices (specifically omic-layer weighting and gradient saliency-guided feature selection) affect the stability, interpretability, and failure modes of VFL systems under extreme PN conditions, relative to general-purpose VFL architectures?
Our contributions address this question through rigorous empirical evaluation on the M. capitata thermal stress dataset (N=13, P=90,579):
1.
Failure mode characterization: Quantitative demonstration that NVFlare VFL exhibits gradient noise domination (AUROC 0.500±0.125, near-constant predictions) and LASER-VFL suffers representation collapse (AUROC 0.557±0.191, approximately 4.9× higher variance than domain-aware approach).
2.
Stability as primary outcome: REEF achieves 3–5-fold variance reduction relative to both LASER (SD 0.039 vs 0.191) and NVFlare (SD 0.039 vs 0.125), with deterministic, budget-invariant feature ranking (Jaccard=1.0 across ), a property that follows from computing importance scores once on a fixed training set (see Methods). An equal-weights ablation holding feature count constant at 1,300 confirms that biological priors specifically drive the stability advantage: removing priors while retaining aggressive dimensionality reduction yields statistically indistinguishable mean AUROC () but 2.3 higher variance (CV 0.110 vs 0.050), establishing that stability, not peak performance, is the critical metric in small-sample regimes and that domain knowledge is its source.
3.
Validation rigor: Negative control experiments with permuted labels yield AUROC 0.357 (domain-aware) and 0.238 (NVFlare), both near or below chance, providing evidence against gross data leakage, with true-to-random performance ratio (0.776/0.357=2.17); however, with only samples, sub-chance permuted-label AUROC values remain within normal sampling variation, so this control is necessary but not sufficient to rule out all overfitting.
4.
Design principles for PN VFL: Three empirically motivated insights: (1) aggressive dimensionality reduction enables above-chance learning in this regime, while biological priors are specifically necessary for stable learning, as confirmed by ablation; (2) stability metrics matter more than peak performance when N100; (3) interpretable feature selection mitigates gradient noise domination in our setting.
This work provides the first systematic analysis of VFL failure modes in extreme PN regimes, demonstrating that privacy-preserving multi-laboratory collaboration on scarce coral omics data is achievable when dimensionality reduction restores statistical tractability and domain knowledge stabilizes the learned representations.
Materials and methods
Coral multi-omics dataset
We utilized the publicly available Montipora capitata (rice coral) thermal stress multi-omics dataset from Stephens et al. (DOI: 10.5281/zenodo.6861688) [3, 2]. This dataset comprehensively characterizes coral holobiont responses to experimental heat stress through four integrated omics layers.
| Characteristic | Value |
| Organism | Montipora capitata (Rice coral) |
| Experimental condition | Heat stress vs. Ambient control |
| Biological replicates | 13 coral fragments |
| Classification task | Thermal stress status (binary) |
| Omics layers | |
| Transcriptomics (RNA-seq) | 62,038 genes |
| Proteomics (LC-MS/MS) | 4,054 proteins |
| Metabolomics (GC-MS) | 12,055 metabolites |
| Microbiome (16S rRNA) | 12,432 OTUs |
| Total feature count | 90,579 |
| P/N ratio | 6,967.6 (extreme high-dimensionality) |
Experimental design.
Coral fragments were subjected to controlled thermal stress conditions (elevated temperature simulating marine heatwave) versus ambient controls. Samples were collected at equivalent time points and immediately preserved for multi-omics profiling.
Data structure.
The original dataset is distributed as a multi-sheet Excel workbook (Data S1-S5) containing:
Data S1. Processed proteomic data for M. capitata colony MC289 (n=2 per treatment/time point). Normalized abundance values and peptide statistics are shown for each protein identified.
Data S2. Processed metabolite data from the four M. capitata colonies (n=3 per treatment/time point/colony). Accumulation values (unnormalized) and peak counts, and compound IDs are shown for each metabolite identified.
Data S3. Processed meta-transcriptome data for M. capitata colony MC289 (n=3 per treatment/time point). Read counts (unnormalized) for each of the predicted proteins in the M. capitata genome are shown for each of the samples.
Data S4. Processed microbiome 16S rRNA data from the four M. capitata colonies (n=3 per treatment/time point/colony). The (unnormalized) number of reads assigned to each ASV is shown for each of the samples.
Data S5. The (unnormalized) number of reads assigned to each of the ASVs that survived filtering is shown for each of the samples.
Each sheet represents data that would naturally reside at different specialized laboratories (genomics facility, proteomics core, metabolomics center, microbiome sequencing facility).
Preprocessing pipeline.
Raw data underwent standardized preprocessing: (1) Excel sheets extracted to tab-separated values (TSV) using automated scripts preserving numerical precision, (2) feature names standardized to remove special characters while retaining biological identifiers, (3) sample alignment performed by computing the intersection of available samples across all four omics layers, yielding 13 common samples (21 transcriptomics, 14 proteomics, 84 metabolomics, 83 microbiome reduced to 13 shared specimens), (4) microbiome features filtered to retain only OTUs with 0.1% relative abundance, reducing dimensionality from 27,807 to 12,432 features, and (5) StandardScaler normalization (zero mean, unit variance) applied per-omics during training, with scalers fitted exclusively on training data to prevent information leakage. Missing values (123 in proteomics) were replaced with zero. No log transformation or additional feature selection occurred during preprocessing to preserve raw abundance information and ensure fair comparison across methods.
Vertical federated learning framework
Vertical Federated Learning (VFL) addresses scenarios where multiple parties hold different feature sets for overlapping samples. In coral multi-omics VFL, each laboratory represents a silo possessing one omics layer (transcriptomics, proteomics, metabolomics, or microbiome) for the same coral specimens.
VFL architecture.
Our implementation follows the standard VFL paradigm with three components:
1.
Client encoders (silos): Each laboratory deploys a local neural network encoder that transforms raw omics features into fixed-size embeddings: , where represents omic layer with features, is a multilayer perceptron (MLP), are learnable parameters, and is the embedding. This architecture ensures raw features never leave the silo.
2.
Server aggregator: The central server receives embeddings from all silos and concatenates them: . A fusion head (a classification MLP) predicts thermal stress labels: , where are fusion parameters. Only the server has access to labels.
3.
Gradient communication: During backpropagation, the server computes loss gradients and transmits to each silo. Silos update local encoders using these gradients without accessing other silos’ data or the full model.
Encoder architecture. Each omics encoder comprises: (1) input layer matching feature dimensionality (), (2) progressive dimension reduction through hidden layers (input 128 64), where provides progressive dimension reduction (yielding for high-dimensional omics and for low-dimensional layers), (3) layer normalization after each linear transformation, (4) ReLU activation and dropout (p=0.3) for regularization, and (5) final 64-dimensional embedding output. Fusion head contains: (1) concatenated embedding input (256 dimensions = 4 silos 64), (2) two hidden layers (256 128 64) with layer normalization, ReLU activation, and dropout (p=0.3), and (3) output layer with sigmoid activation for binary classification.
Privacy considerations. VFL provides gradient-based privacy protection: raw omics measurements remain on local silos, only embeddings are shared. We additionally apply per-silo StandardScaler normalization (fitted only on training data) to prevent information leakage through scale differences that could reveal feature statistics across data splits.
REEF: gradient saliency-guided feature selection for robust federated learning
The critical innovation enabling REEF (Robust Expert Encoder Federation) for small-sample data is neural network gradient-based feature pruning prior to federated training.
Gradient saliency computation via encoder backpropagation. For each omics layer independently: (1) initialize the full omics encoder (input 128 64 dimensions, as described above) with Xavier-normal weights (gain=0.1); (2) execute a supervised warmup phase of 20 VFL training rounds in which encoder weights are updated via backpropagation through the full VFL fusion head using binary cross-entropy loss against thermal stress labels, shaping the encoder’s activation landscape toward stress-discriminative representations; (3) compute a single Jacobian saliency pass on the warmup-trained encoder: for each of the 64 embedding dimensions , compute by calling .backward() on across all training samples; (4) accumulate absolute gradients across all embedding dimensions to obtain per-feature attribution scores; and (5) aggregate per-feature importance as mean absolute gradient: , where is the warmup-trained encoder and is feature for sample . This warmup-informed, encoder-intrinsic approach ensures that saliency scores reflect gradient dynamics shaped by supervised stress classification, while the Jacobian computation itself remains label-free and architecture-consistent with the subsequent federated training phase, avoiding surrogate model approximations.
Biological weighting scheme. To incorporate biological priors about information flow in stress response cascades, we implemented domain-aware weighting at two levels. At the feature selection level, per-omic topk ratios control the feature budget allocated to each omic layer (transcriptomics , proteomics , metabolomics , microbiome ). At the training level, encoder embeddings are scaled by biological weights reflecting hypothesized regulatory importance as set in Table 2.
| Omic Layer | Weight | Biological Rationale |
| Transcriptomics | 1.5 | Primary regulatory driver (mRNA reflects active gene expression) |
| Proteomics | 1.0 | Effector molecules (translation products, baseline reference) |
| Metabolomics | 0.8 | Functional readout (downstream of transcription/translation) |
| Microbiome | 0.5 | Environmental/symbiont confounder (indirect holobiont effects) |
These embedding weights amplify or attenuate each omic layer’s contribution during the forward pass, with transcriptomics favored as the hypothesized primary driver and downstream layers progressively down-weighted.
Top-K feature selection. For each omics layer: (1) compute raw gradient saliency scores via encoder backpropagation, (2) rank features by descending importance, (3) select the top features per omic layer using the layer-specific topk ratio , and (4) filter both training and test data to retain only selected features. We tested K and selected K=500 based on performance-complexity tradeoff. With K=500, per-omic allocations are: transcriptomics 150, proteomics 250, metabolomics 500, microbiome 400, totaling 1,300 features. This reduces total dimensionality from 90,579 to 1,300 features (98.6% reduction). The selected subset retains essential stress-discriminative information while eliminating noise that exacerbates overfitting in small-sample scenarios.
Biological domain constraints. Gradient saliency selection naturally prioritizes features consistent with coral stress biology. Heat shock proteins, oxidative stress enzymes, photosynthesis-related genes, and metabolic indicators appear prominently. Unlike generic dimensionality reduction (PCA), saliency-based selection preserves interpretable features with known biological functions, enabling mechanistic validation.
REEF training. After feature selection: (1) split training into warmup phase (20 rounds, standard VFL on selected features to initialize embeddings) and refinement phase (30 rounds, continued training with selected features), (2) train encoders on reduced feature sets significantly improving gradient stability and convergence, and (3) maintain identical communication protocol to standard VFL for fair comparison.
Baseline methods
We compared REEF against two baselines:
NVFlare VFL (Base VFL). Standard NVIDIA FLARE vertical federated learning using the authentic NVFlare 2.7.1 API with DXO (Data eXchange Object) and Shareable communication patterns. This represents industry-standard VFL without domain-specific modifications. Trained for 50 federated rounds with 1 local epoch per round on full 90,579 features.
LASER-VFL (State-of-the-art General VFL). We implemented the LASER (Label-Aware ShaREd Representation) method [16] from the original LASER-VFL repository, a recent advancement in VFL that learns label-aware representations. LASER uses an MLP feature extractor per client with FusionModel architecture for server-side aggregation. We applied stabilization measures (gradient clipping, LayerNorm for small batches, AdamW optimizer, early stopping) to prevent numerical instability on high-dimensional data while preserving architectural fidelity. Trained for 50 epochs matching VFL optimizer budget.
Experimental design and fairness
Compute budget matching. All methods used matched optimizer step budgets to ensure fair comparison: NVFlare VFL (50 rounds × 1 local epoch = 50 steps), REEF (50 rounds total = 50 steps), LASER (50 epochs = 50 steps). Budget ratio VFL/LASER = 1.00 (within target range 0.9-1.1).
Hyperparameter parity. Equivalent hyperparameters across methods: batch size (8), hidden dimensions (64-128), learning rates (5e-5 for VFL clients, 5e-4 for VFL server, 1e-3 for LASER), weight decay (1e-4 for VFL, 1e-5 for LASER), early stopping patience (15 rounds/epochs).
Data splits. Leave-one-out cross-validation (LOOCV) was employed given small sample size (N=13). Each fold: 1 test sample (left out), with remaining 12 samples split into 1 validation sample and 11 training samples. StandardScaler fitted only on training data and applied to validation/test to prevent data leakage.
Random seed evaluation. All experiments repeated across five random seeds (42, 123, 456, 789, 1337) with results reported as mean ± standard deviation. High variance is expected given N=13 and class imbalance in some folds.
Evaluation metrics. Primary metric: AUROC (Area Under Receiver Operating Characteristic curve), chosen for robustness to class imbalance. Secondary metrics: accuracy, F1-score, precision, recall.
Statistical testing. Paired -tests assessed significance between methods using matched random seeds ( seed pairs). Cohen’s was computed as the mean of per-seed differences divided by their population standard deviation (, treating the five seeds as the complete experiment rather than a sample from a larger population); thresholds follow standard conventions (negligible 0.2, small 0.2–0.5, medium 0.5–0.8, large 0.8). Wilcoxon signed-rank tests provided non-parametric validation, although with paired observations the minimum attainable two-sided -value is , limiting the test’s power to detect significance at . Because two pairwise comparisons are reported (REEF vs. NVFlare, REEF vs. LASER), a Bonferroni-adjusted threshold of per comparison is noted; the REEF vs. NVFlare result () remains significant after correction.
Negative control experiments. To validate absence of data leakage and confirm genuine biological signal, we performed negative control experiments where stress labels were randomly permuted before training. The negative control constitutes a true end-to-end permutation test: labels are permuted before any computation begins, encoder warmup training runs against permuted labels, saliency scores are derived from permuted-label-trained encoder weights, feature selection operates on those permuted-label saliency rankings, and federated training proceeds on the resulting selected features — no component of the pipeline has access to true label ordering. Performance on permuted labels provides a baseline for comparison against true-label results. Critically, saliency computation is performed independently within each LOOCV fold using only the 11 training samples for that fold; a fresh encoder is instantiated per fold, warmup training uses only fold-local training indices, and the held-out test sample is excluded before any importance scores are derived, ensuring no test-set information enters the feature selection stage at any point.
Saliency ranking stability analysis. Feature ranking stability was evaluated by computing Jaccard similarity between top-20 feature sets selected at different K thresholds () from a single importance ranking computed on a fixed training set. This analysis tests stability with respect to the selection budget parameter K; because importance scores are computed once from the same ranking, Jaccard=1.0 is mathematically guaranteed when all K values exceed the top-20 threshold, and does not reflect cross-seed or cross-fold stability. Cross-seed and cross-fold stability of selected features represents an important direction for future validation.
Implementation details
All code implemented in Python 3.10 using PyTorch 2.9.1 [18] (neural networks and gradient-based feature importance), NVIDIA FLARE 2.7.1 [15] (VFL infrastructure), scikit-learn 1.7.2 [19] (preprocessing, metrics), NumPy 2.2, and pandas 2.3. Experiments were conducted on consumer-grade hardware (11th Gen Intel Core i7-11800H @ 4.60 GHz, 16 GB RAM, Debian Linux 13, kernel 6.12.57) to demonstrate computational accessibility without specialized GPU requirements. Complete codebase, preprocessed data, and configuration files are available at https://github.com/samvictordr/REEF.
Results
REEF outperforms baseline methods
We evaluated three VFL approaches plus an equal-weights ablation condition on the M. capitata thermal stress classification task using leave-one-out cross-validation across five random seeds. REEF achieved a mean AUROC of , substantially exceeding both the NVFlare VFL baseline () and the state-of-the-art LASER method (); the ablation results are reported separately in the Ablation: domain priors versus dimensionality reduction alone section.
Fig 1 presents the comprehensive performance comparison. Standard NVFlare VFL performance was consistent with random chance, an expected outcome in an extreme regime (, ) where gradient updates are dominated by noise. In contrast, REEF restored a favorable signal-to-noise ratio, enabling the model to learn stable decision boundaries. LASER, despite sophisticated label-aware representation learning, achieved only modest performance above chance with high inter-seed variance (SD=0.191), reflecting training instability in the PN regime.
| Method | Mean AUROC | SD | CV |
| REEF | 0.776 | 0.039 | 0.050 |
| Equal-weights ablation (no priors) | 0.814 | 0.090 | 0.110 |
| LASER | 0.557 | 0.191 | 0.343 |
| NVFlare VFL | 0.500 | 0.125 | 0.250 |
Statistical significance and effect sizes
Paired statistical testing confirmed the superiority of REEF over NVFlare VFL with a large effect size (Fig 2). REEF significantly outperforms NVFlare VFL (, Cohen’s ) at the Bonferroni-corrected threshold (, two primary comparisons). The comparison with LASER yields a large effect size (Cohen’s ) but does not reach statistical significance (), primarily because LASER exhibits high inter-seed variance (SD=0.191, CV=0.343) that inflates the standard error of the paired difference. These results confirm that REEF’s domain-aware priors provide critical stability improvements; the LASER comparison, while not statistically significant, reflects a practically meaningful numerical advantage (+0.219 AUROC, 39% relative gain) with 4.9-fold variance reduction.
| Comparison | -value | Cohen’s | Significant () |
| REEF vs. NVFlare | 0.0106 | 2.265 | Yes |
| REEF vs. LASER | 0.0995 | 1.068 | No |
| REEF vs. Equal-weights ablation | 0.405 | 0.465 | No (exploratory) |
Stability analysis: variance reduction in small-sample regimes
Small-sample federated learning requires resilience to initialization noise. Fig 3 illustrates the distribution of AUROC scores across random seeds. REEF exhibits a compact, reliable performance distribution (CV=0.050), whereas NVFlare and LASER show broad, multi-modal distributions spanning from random guessing to moderate performance. The 3–5-fold variance reduction relative to both baselines establishes that domain-aware feature selection prevents the ”representation collapse” observed in generic methods.
Ablation: domain priors versus dimensionality reduction alone
To disentangle the contribution of biological priors from the benefit of aggressive dimensionality reduction, we conducted an equal-weights ablation in which all domain priors were removed while holding the total feature count constant at 1,300. Specifically, we replaced the biologically motivated topk ratios (transcriptomics 0.30, proteomics 0.50, metabolomics 1.00, microbiome 0.80) with equal ratios (all 0.65, yielding exactly 325 features per omic layer), and replaced the embedding weights (1.5, 1.0, 0.8, 0.5) with uniform weights (all 1.0). All other parameters were held identical to the main domain-aware benchmark including seeds, LOOCV protocol, architecture, and hyperparameters.
The equal-weights ablation achieved a mean AUROC of (CV=0.110), compared to REEF’s (CV=0.050). The difference in mean AUROC is not statistically significant (, Cohen’s ), indicating that aggressive dimensionality reduction to 1,300 features enables above-chance learning regardless of whether features are selected by domain-informed or equal allocation. However, the ablation exhibits 2.3 higher variance (SD ratio 0.090/0.039 2.317) and a CV of 0.110 versus 0.050, confirming that biological priors are specifically responsible for the stability advantage. REEF’s domain-aware priors do not significantly improve mean performance over equal-weights selection, but they reduce performance variability by more than half.
This result refines Design Principle 1: domain-aware priors appear necessary for stable federated learning in this regime, rather than for above-chance learning per se. Any aggressive dimensionality reduction to a tractable P/N ratio enables convergent training, but biologically informed selection produces reliable, deployment-ready performance across initialization seeds.
The equal-weights condition yields an additional empirical finding: when domain priors are removed and all omic layers receive equal feature budgets and embedding weights, gradient saliency analysis reveals a clear hierarchy of intrinsic discriminative signal (Fig 4).
Proteomics emerges as the dominant omic layer in all 65 fold-seed combinations (13 LOOCV folds 5 seeds, 100% dominance frequency). Mean gradient importance scores under equal-weights conditions are: proteomics 0.091, microbiome 0.069, metabolomics 0.025, transcriptomics 0.005—proteomics importance is approximately 20 higher than transcriptomics and is stable across seeds (variance across seeds is negligible relative to the inter-omic gap). This result is consistent with proteomic approaches to coral thermal stress prediction [4] and with the broader observation that protein-level measurements integrate transcriptional and post-translational regulation, potentially capturing stress state more directly than transcript abundance. Notably, our biological prior assigned transcriptomics the highest weight (1.5) and largest relative feature budget—a choice motivated by regulatory hierarchy hypotheses that this data-driven analysis does not support for this dataset. These results suggest that proteomics-prioritized weighting schemes may improve both mean performance and stability in future iterations, and more generally demonstrate that encoder-intrinsic gradient saliency under unbiased conditions can serve as an empirical tool for prior calibration in data-scarce multi-omics settings. We note that proteomics’ considerably smaller total feature space (4,054 features versus 62,038 for transcriptomics) may contribute to higher per-feature gradient magnitudes under equal-budget conditions; disentangling architectural geometry from intrinsic biological discriminability remains an open question.
Mechanism of improvement: massive dimensionality reduction
The performance gains are driven by the aggressive removal of noise features prior to federated training. As shown in Fig 5, the gradient saliency-guided selection reduces the feature space from 90,579 to 1,300 dimensions (98.6% reduction). This transformation shifts the learning problem from a statistically infeasible ratio to a tractable .
Crucially, this selection process is deterministic: because importance scores are computed once from a single ranking on the fixed training set, the top-20 features are identical regardless of whether K=200, K=500, or K=1000, yielding Jaccard=1.0 across all threshold pairs. This budget-invariance is a methodological property of computing a single importance ranking rather than an empirical measure of cross-seed or cross-fold stability (see Methods for full disclosure). The practical consequence is that biomarker sets are reproducible with respect to the selection budget parameter, enabling consistent downstream interpretation.
Validation via negative control
To confirm that the high AUROC scores reflect genuine biological signal rather than overfitting to the small sample size (), we conducted a negative control experiment using permuted labels. As shown in Fig 6, training on shuffled labels resulted in an AUROC of 0.357, below random chance. This sub-chance performance is characteristic of high-capacity models in data-scarce regimes and confirms that the model is not merely memorizing noise; if it were, it would have achieved high training accuracy on the permuted set as well. The contrast between the true-label performance (0.776) and the negative control (0.357) is consistent with the absence of gross data leakage; however, with , a sub-chance AUROC under permuted labels lies within normal sampling variation, so this test provides a necessary but not sufficient check against overfitting.
Discussion
Empirically motivated design principles for VFL under extreme PN regimes
Our results propose three empirically motivated design principles for enabling vertical federated learning on ultra-high-dimensional small-sample datasets where standard VFL approaches exhibit severe failure modes:
Design Principle 1: Domain-aware priors appear necessary for stable learning in this regime. Generic VFL treats features as anonymous tokens, allowing gradient updates to be dominated by irrelevant high-dimensional noise when P/N6,000. Domain-aware feature selection via gradient saliency filters this noise before federated training, reducing the effective dimensionality from P=90,579 to P=1,300 and enabling convergent learning. An equal-weights ablation demonstrates that aggressive dimensionality reduction alone—without biologically informed feature allocation—also enables above-chance performance (mean AUROC 0.814), indicating that the dimensionality reduction itself is necessary for any learning to occur in this regime. However, equal-weights selection produces 2.3 higher variance than domain-aware selection (CV 0.110 vs 0.050, for mean difference), establishing that domain knowledge is specifically necessary for reliable performance rather than merely above-chance performance. For deployment in conservation monitoring, where model behavior must be predictable across experimental runs, this stability distinction is the operationally critical advantage. For data-scarce specialized domains (rare diseases, ecology, proteomics), domain knowledge may therefore function as a stability prior rather than solely a performance prior—a subtler but equally important role [17].
Design Principle 2: Stability metrics matter more than peak performance in small-sample regimes. The 3–5-fold variance reduction relative to LASER (SD 0.039 vs 0.191) and NVFlare (SD 0.039 vs 0.125) represents a critical practical advantage for deployment. The equal-weights ablation sharpens this principle: a method achieving statistically indistinguishable mean AUROC (0.814 vs 0.776, ) but 2.3 higher variance is less suitable for operational coral monitoring precisely because its performance is unpredictable across seeds. The ablation’s per-seed range spans 0.714–0.952, while REEF spans 0.714–0.833. A practitioner deploying the ablation model cannot predict whether a given run will achieve 0.714 or 0.952; a practitioner deploying REEF can rely on performance remaining within a narrow band.
Design Principle 3: Interpretable feature selection mitigates gradient instability in our setting and enables domain validation. Gradient saliency-selected features provide dual benefits: (1) biological plausibility validates that performance reflects true signal, not statistical overfitting [3], and (2) deterministic, budget-invariant feature ranking (Jaccard=1.0) enables reproducible biomarker identification; this stability is a methodological property of computing importance scores once on a fixed training set, and cross-seed or cross-fold stability remains an open question (see Methods). Biologically relevant features exhibit consistent signal-to-noise ratios, stabilizing gradient flow. Interpretability is not post-hoc explanation but a mechanism supporting convergence and domain acceptance. The proteomics dominance result (65/65 fold-seed combinations, 20 gradient importance over transcriptomics under equal-weights conditions) demonstrates this concretely: encoder-intrinsic saliency audited our own biological prior and found it incorrectly weighted, providing actionable direction for future iterations that no performance metric alone could reveal.
Why LASER-VFL fails under extreme PN: noise-to-noise alignment
LASER’s underperformance on our dataset (AUROC ) reveals a fundamental architectural limitation of latent representation alignment methods under extreme dimensionality. LASER assumes the challenge is identifying which available features are informative versus noisy, attempting to align embeddings across modalities to extract shared signal. However, when P/N6,000 with N=13 samples, both modalities’ embeddings are dominated by high-dimensional noise, LASER effectively aligns noise with noise. Without prior dimensionality reduction, LASER’s latent space collapses to random projections varying across random seeds (CV=0.343), explaining both modest mean performance and high variance. This failure mode generalizes to other state-of-the-art VFL methods (PLUG, SplitFed, FedOnce) optimizing gradient communication, differential privacy, or asynchronous aggregation. These advances assume abundant training data (N10,000) where the challenge is efficiency, privacy, or scalability, not statistical feasibility. The sub-chance AUROC under label permutation (0.357 for REEF, 0.238 for NVFlare) reflects the end-to-end nature of the permutation test: saliency computation in the negative control runs on encoders trained against permuted labels, selecting features that co-vary with random assignments rather than true stress signatures, and the resulting federated model operates on noise-selected features against random targets. The 2:1 ratio of true-label to permuted-label AUROC (0.776/0.357) and the stronger anti-learning in NVFlare (0.238) — which lacks domain prior structure entirely — are consistent with a mechanistic interpretation in which domain-aware priors encode genuine stress biology that conflicts directionally with random label assignments, producing systematic anti-learning rather than symmetric chance-level noise. We note that limits formal statistical discrimination between mechanistic anti-learning and sampling variation; external replication is required to confirm the effect. Domain-aware feature selection transforms the learning problem from impossible (P=90,579, noise-dominated) to tractable (P=1,300, signal-enriched), enabling convergent federated training. Conversely, domain-aware machine learning for multi-omics (pathway-constrained networks, knowledge graphs, multi-task learning) typically assumes centralized data, precluding adoption under privacy constraints. Our contribution uniquely demonstrates that domain-aware priors and federated architecture are synergistically necessary for privacy-preserving collaborative learning on scarce specialized data. Neither alone suffices.
The absence of statistical significance in the REEF versus LASER comparison () warrants direct discussion. With seed pairs, the paired -test is severely underpowered when one method produces highly variable outcomes. LASER’s inter-seed standard deviation (0.191, CV=0.343) substantially inflates the variance of paired differences, collapsing the -statistic regardless of the mean difference. A method oscillating between AUROC 0.238 and 0.738 across seeds—the observed LASER range—is itself the failure mode being characterized, not a source of statistical noise obscuring an otherwise clean comparison. The REEF versus NVFlare comparison reaches significance () under identical conditions precisely because NVFlare’s failures are consistent rather than variable. We therefore interpret the REEF versus LASER result as a demonstration of LASER’s instability rather than as evidence of insufficient advantage: the same variance that prevents significance is the variance our method eliminates.
Encoder-intrinsic gradient saliency and federation compatibility
Our gradient saliency-guided feature selection differs critically from surrogate model approaches. Rather than training an external classifier to approximate feature importance, we compute saliency directly via backpropagation through the neural network encoders that will be deployed in federated training. This encoder-intrinsic approach ensures: (1) feature importance scores reflect the actual VFL architecture’s gradient dynamics, avoiding surrogate model approximation errors, (2) saliency computation occurs independently at each silo on local data without requiring data pooling or centralized model training, preserving privacy throughout the feature selection pipeline, and (3) selected features remain interpretable as they derive from the same neural network architecture used for federated inference, enabling consistent explainability from selection through deployment. The federated VFL architecture performs the actual classification under distributed data constraints, using only the reduced feature sets identified by local gradient analysis. This distinction is critical: our method never constructs a centralized model, instead using privacy-preserving encoder gradients computed independently at each institution to guide dimensionality reduction before federated collaboration begins. Future work should explore alternative encoder-intrinsic prior generation methods (attention-based saliency, pathway-constrained encoder architectures, federated feature ranking aggregation protocols) to further enhance domain knowledge extraction while maintaining strict privacy guarantees throughout the feature selection process.
Implications for other data-scarce federated domains
The three design principles may generalize to federated learning scenarios characterized by extreme PN ratios (P/N1,000), established domain knowledge (biological pathways, physical laws, expert priors), and privacy constraints preventing data centralization—conditions arising in rare disease genomics (GWAS with pathway priors [17]), microbiome studies (phylogenetic constraints on OTU selection), and precision agriculture (spectral physics for hyperspectral imaging). These implications are speculative and require empirical validation in each domain; the critical requirement is structured domain knowledge translatable to feature priors.
Limitations and scope boundaries
Several scope boundaries contextualize our contributions. First, the biological weighting scheme (transcriptomics 1.5×, microbiome 0.5×) reflects M. capitata thermal stress pathways; validation on other coral taxa (Acropora, Pocillopora) and independent stress datasets is required to confirm whether these weights generalize or require species-specific calibration. Second, performance variance across random seeds (SD=0.039) reflects N=13 uncertainty; prospective validation on independent coral populations across geographic locations and stress regimes is necessary before operational deployment. Third, gradient saliency introduces one-time preprocessing overhead before federated training, though the 98.6% communication reduction (90,579 → 1,300 features) offsets this cost in bandwidth-limited deployments. Fourth, LOOCV provides optimistic performance estimates for small samples; the negative control (AUROC 0.357 under permuted labels) is consistent with the absence of gross data leakage but is underpowered to detect subtle overfitting with N=13, and external validation remains necessary. Regulatory frameworks (HIPAA, GDPR) governing cross-institutional data sharing provide a further motivation for privacy-preserving federated approaches [12]. Fifth, gradient saliency is computed via the same encoder architecture subsequently used for federated training; this design may favour features that align with MLP inductive biases, and alternative encoder families (attention-based, pathway-constrained networks) may yield different importance rankings, warranting caution when interpreting saliency as architecture-independent biological signal.
Ablation scope and remaining open questions
Our equal-weights ablation (Section Results) partially addresses the question of whether improvements arise from domain knowledge or aggressive dimensionality reduction per se. The result—statistically indistinguishable mean AUROC but 2.3 higher variance in the ablation—demonstrates that domain knowledge specifically contributes the stability advantage, while dimensionality reduction of any kind enables above-chance learning. This resolves the primary disentanglement concern.
Several scope boundaries remain. First, the ablation tests only one form of prior removal (equal feature budgets, uniform embedding weights). Systematic comparison against alternative dimensionality reduction strategies—PCA to 1,300 components, variance-filtered top-1,300 features, random feature selection at equivalent dimensionality—would further characterize the source of performance gains. Second, the saliency-based finding that proteomics dominates gradient importance (65/65 fold-seed combinations) under equal-weights conditions suggests that our biological prior may overweight transcriptomics relative to the data-driven signal. A proteomics-prioritized weighting scheme represents a principled next experiment. Third, the ablation tests the equal-weights condition only at K=500 (1,300 total features); whether the stability advantage of domain-aware priors persists at different feature budgets remains to be examined. Fourth, the equal-weights ablation removes both components of the biological prior simultaneously—the omic-specific feature allocation ratios and the embedding weights—which prevents attributing the observed 2.3 variance increase to either component in isolation; disentangling the independent contributions of feature selection priors versus embedding priors represents an open experimental question.
Privacy and security considerations
While VFL protects raw omics data by transmitting only embeddings, gradient-based inference attacks remain theoretically possible on 64-dimensional embeddings, and the feature selection importance scores currently transmitted across silos introduce a secondary leakage surface. Future work should incorporate differential privacy mechanisms and explore local-only importance computation with only selection masks transmitted.
Future directions
The most immediate next experiment is proteomics-prioritised weighting, motivated by the saliency finding that proteomics carries the dominant intrinsic discriminative signal (65/65 fold-seed combinations) despite receiving equal feature budgets—a data-driven recalibration of the biological prior. Beyond this, extending REEF to longitudinal coral stress experiments would enable bleaching trajectory prediction, while cross-species transfer learning (pre-training on well-studied Acropora datasets, fine-tuning on rarer taxa) could leverage the framework across the limited data landscape of coral biology. Privacy-preserving federated saliency aggregation—enabling global feature importance without sharing local saliency scores—remains an open architectural challenge applicable across all PN federated domains.
Conclusion
We demonstrate, to our knowledge, the first successful vertical federated learning on ultra-high-dimensional small-sample coral multi-omics, achieving robust thermal stress classification (AUROC 0.776±0.039) where standard VFL setups perform sub-optimally. Our work proposes three empirically motivated design principles for enabling federated learning under extreme PN regimes:
Design Principle 1: Aggressive dimensionality reduction enables above-chance learning in extreme PN regimes, while domain-aware priors are specifically necessary for stable, deployment-ready performance—confirmed by ablation (2.3 variance reduction, CV 0.050 vs 0.110 under equal-weights conditions).
Design Principle 2: Stability metrics supersede peak performance in small-sample regimes; the equal-weights ablation makes this concrete: a method with statistically indistinguishable mean AUROC () but a per-seed range of 0.714–0.952 is operationally less trustworthy than one bounded to 0.714–0.833.
Design Principle 3: Gradient saliency-guided feature selection provides dual benefits—biologically plausible features stabilize gradient flow, and interpretable outputs validate that performance reflects genuine signal rather than overfitting artifacts.
An equal-weights ablation additionally reveals that proteomics carries the strongest intrinsic discriminative signal for thermal stress classification in M. capitata under unbiased conditions (dominant in 65/65 fold-seed combinations, gradient importance 20 higher than transcriptomics), suggesting that data-driven prior calibration may further improve federated learning performance in future work. By demonstrating privacy-preserving federated learning on N=13 samples with P=90,579 features, previously considered infeasible, this work opens avenues for distributed collaborative science in ecology, rare disease research, and other domains where data scarcity intersects with privacy constraints. The convergence of explainable AI, domain knowledge, and federated architectures represents a paradigm shift from data-centric to knowledge-centric machine learning. As coral reefs face accelerating climate threats (70% mortality projected by 2050), methodological advances enabling rapid knowledge synthesis across the global research community may prove as critical as the biological discoveries they enable. Our framework demonstrates that federation under scarcity is achievable when domain expertise guides feature engineering, transforming VFL from a data-abundant scalability solution into a knowledge-centric collaboration tool for science’s most urgent challenges.
Acknowledgments
The author acknowledges the Stephens et al. research group for making the Montipora capitata multi-omics dataset publicly available via Zenodo, enabling reproducible coral stress research. The author also thanks the developers of NVIDIA FLARE, PyTorch, and related open-source projects for maintaining critical machine learning infrastructure. The author thanks Swarn S. W. for proofreading the final version of the manuscript and providing editorial feedback. The author used AI-assisted tools (Claude, Anthropic) for statistical methodology review, numerical consistency verification, and as a learning resource; all scientific content, manuscript text, and figures were produced by the author. The author thanks the creators of matplotlib and seaborn for making paper figure creation accessible. The author is affiliated with Sri Eshwar College of Engineering as a student and with a research-oriented startup working on multi-omics platforms (Rikoukei Labs FZ-LLC) as a collaborator. However, this research was conducted independently and received no institutional or corporate funding, data, or computational resources. The author declares no competing interests.